OpenStudy (anonymous):

whats the square root of -49 ?

OpenStudy (stefrheart):

DO you know about imaginary numbers?

OpenStudy (anonymous):

7i

OpenStudy (traceur55):

The ninja by samliedoops...

OpenStudy (anonymous):

haha what?!

OpenStudy (anonymous):

Do you need me to show work ?

OpenStudy (traceur55):

A "ninja" is replying with the answer right before someone else does. I was literally about to click post. :p

OpenStudy (anonymous):

\[\sqrt{-49}\] \[\sqrt{(-1)(49)}\] \[\sqrt{-1}\sqrt{49}\] = i * 7 i7

OpenStudy (anonymous):

Haha, I'm sorry!

OpenStudy (anonymous):

huh ?